OpenStudy (anonymous):

find the coordinates of the focus, of the parabola y=x^2+4 !! PLEASE HELP

OpenStudy (anonymous):

standard form: y = x^2, 1/(4p) =1, so p = 1/4 The parabola opens up, focus is (0,1/4) y=x^2 +1, y-1 = x^2, you decrease 1 value of y, so need to move up 1 unit of y-coordinate to maintain equality. (0,1/4) ----shift-----focus: (0,1/4+1)

OpenStudy (anonymous):

the equation actually was y=x^2+4

OpenStudy (anonymous):

so would the focus just be (0,1+4)

OpenStudy (anonymous):

it's (0,1/4+4)

OpenStudy (anonymous):

so for this equation y=x^2+4 the focus is (0,14+4) ?

OpenStudy (anonymous):

(0, 17/4). 1/4 is fraction

OpenStudy (anonymous):

i don't get how you got 1/4 ?

OpenStudy (anonymous):

oh, the form of parabola is y=[1/(4p)]x^2 so in this case 1/(4p)=1, so p=1/4

OpenStudy (anonymous):

your confusing me... how did u get that form ?

OpenStudy (anonymous):

it requires a lot of proving. Is this calculus 2 problems?

OpenStudy (anonymous):

its locus

OpenStudy (anonymous):

I never heard of that term. is that the point where distance from the parabola = distance from diretrix?